OpenStudy (anonymous):

PLEASE HELP!!! subtract. 2 16/21 - (-8 5/21). reduce if possible. a)-6 11/21 b)10 11/21 c)11 d)-5

OpenStudy (anonymous):

\[2 \frac{ 16 }{21 }-\left( -8 \frac{ 5 }{ 21 } \right)=2\frac{ 16 }{ 21 }+8\frac{ 5 }{ 21}\] \[2+\frac{ 16 }{ 21 }+8+\frac{ 5 }{ 21 }=10+\frac{ 16+5 }{21 }=10+\frac{ 21 }{ 21 }=10+1=11\]

OpenStudy (anonymous):

or \[2\frac{ 16 }{21 }+8\frac{ 5 }{ 21 }=\frac{ 42+16 }{21 }+\frac{ 168+5 }{ 21 }=\frac{ 58 }{ 21 }+\frac{ 173 }{21}=\frac{ 58+173 }{ 21}=\frac{ 231 }{ 21 }\] =11

OpenStudy (anonymous):

thanks